# Project Page: https://jbruner23.github.io/consistyle/

### v1

@@ 4.4. Ablations
To better understand the impact of key design choices in our technique, we conduct a series of ablation studies, each isolating a critical component to assess its influence on consistency, style alignment, and overall image quality. We first explore the importance of applying AdaIN to subject values, which is critical for preventing style leakage in the attention crossing mechanism. In Fig. 10 we show results when removing the AdaIN operation and using direct Values, as can be seen, omitting this normalization can lead to significant color contamination, especially when anchor images feature strong color biases (e.g., grayscale or highly saturated tones). This effect can cause unintended color shifts across different outputs, undermining the desired stylistic consistency and reducing overall visual coherence. We demonstrate in the last row of Fig. 10 that this issue also occurs when applying AdaIN for subject Values in Consistory, due to their direct hidden state injection causing a similar effect.
We next assess the impact of attention transfer on consistency, evaluating the importance of injecting both keys and queries. We test four configurations: injecting both keys and queries, only keys, only queries, and neither. The results depicted in Fig. 9 indicate that both components contribute to achieving high consistency, as omitting either reduces the generated outputs’ alignment, often leading to artifacts such as the flowing book or color misalignment.
Figure 6. Demonstration of the method’s limitations. The first row illustrates inconsistencies in generating a complex object (spaceship), where high visual detail leads to variation across images. The second row highlights a failure to align with a distinct style—specifically, the Papercraft Collage style, evident in the face details.
[…]
@@ 5. Limitations
Our approach has several limitations, as can be seen in Fig. 6. First, similar to Consistory (Tewel et al., 2024), it can produce suboptimal results when the correspondence mappings or cross-attention masks fail to accurately capture the intended relationships between images. Second, consistency can degrade for subjects with complex structures, such as large vehicles like boats and spaceships, where fine details are often challenging to maintain. This effect can extend to human subjects, where intricate facial features or clothing elements may vary across generations. The phenomenon is more prevalent when the initial subject interpretations differ significantly between images, such as generating a traditional wooden ship in one image and a modern yacht in another for the same “boat” prompt. While this issue can sometimes be mitigated by selecting different seeds, it remains a potential weakness in cases where precise subject alignment is critical.
Finally, highly distinctive artistic styles, such as Papercraft Collage, Voxel Art or other niche 3D aesthetics, can occasionally lead to style misalignment, as these styles often introduce unique structural deformations or exaggerated textures that challenge the statistical alignment technique we employ.
@@ 6. Conclusion
We present Consistyle, a training-free approach for improving consistency while preserving style alignment in text-to-image generation. Our method leverages attention manipulation to enable controlled characteristic sharing between images, even in cases with significant appearance differences. This demonstrates that consistent image generation is feasible in style-diverse contexts, despite the typical entanglement of style and content that often challenges conventional approaches.
Appendix A User Study Instructions and Questions

### v2

@@ 4.5. Ablations
To better understand the impact of key design choices in our technique, we conduct a series of ablation studies, each isolating a critical component to assess its influence on consistency, style alignment, and overall image quality. The studied variants are:
  1. (i)
    Consistory with an increased step budget (75 steps) to equate the runtime to our method,
  2. (ii)
    Consistyle without query injection,
  3. (iii)
    Consistyle without key injection,
  4. (iv)
    Consistyle without both query and key injection,
  5. (v)
    Consistyle with direct identity injection between images instead of subject-based DIFT mapping,
  6. (vi)
    Consistyle without AdaIN in the attention crossing.
The quantitative outcomes are summarized in Tab. 5, with representative examples shown in Fig. 9 and Fig. 10. Variant (i) demonstrates that extending Consistory’s runtime does not substantially change its performance: our method continues to surpass it across all metrics except DreamSim, consistent with earlier results. Variants (ii)–(v) confirm the importance of the attention injection design. Removing the query, key, or both worsens similarity scores, since the attention injection mechanism is designed to enhance subject consistency. At the same time, its absence slightly improves some style-alignment metrics, as fewer details are transferred across images.
In variant (v), we replace the DIFT-based mapping with a full direct identity injection from an anchor image. Since no patch-based mapping is available and subject segmentations vary in size, we inject keys and values from the entire image. In addition, since no distance measures are available to compare across different anchors, the injection is restricted to the first image in the batch. As shown in Tab. 5 and illustrated in Fig. 9, this variant leads to feature leakage across non-subject regions, resulting in degraded style-alignment scores. In particular, the image backgrounds shift noticeably in their tone compared to their original SDXL counterparts, unlike other variants where the backgrounds remain nearly unchanged.
Finally, variant (vi) highlights the role of AdaIN: without it, values are transferred directly across images, which weakens style alignment, as reflected by a pronounced drop in the Gram metric and visible artifacts in Fig. 10.
Overall, all ablated versions perform worse than our full model, underscoring the necessity of each component in achieving both strong consistency and style fidelity.
Table 5. Comparison results of the ablation study. Bold values denote the best results and underlined values indicate the second best.(Mean ± SD)
Ablation Method DreamSim $\downarrow$ CLIPScore $\uparrow$ CLIPScore, Styled $\uparrow$ LPIPS $\downarrow$ Gram L2 $\downarrow$ DINO $\uparrow$
Consistory (original, 50 steps) $\mathbf{0.33\pm 0.12}$ $32.75\pm 1.53$ $35.34\pm 1.67$ $0.27\pm 0.07$ $1.85\pm 1.13$ $0.78\pm 0.11$
(i) Consistory (75 steps) $0.40\pm 0.12$ $32.79\pm 1.55$ $35.40\pm 1.67$ $0.42\pm 0.04$ $1.57\pm 0.30$ $0.77\pm 0.11$
Consistyle (full) $\underline{0.40\pm 0.10}$ $\underline{32.84\pm 1.66}$ $\mathbf{36.03\pm 1.69}$ $0.21\pm 0.06$ $1.25\pm 0.69$ $0.85\pm 0.08$
(ii) Consistyle (no Q injection) $0.48\pm 0.11$ $32.75\pm 1.64$ $\underline{35.99\pm 1.61}$ $0.28\pm 0.03$ $\underline{1.18\pm 0.23}$ $\underline{0.87\pm 0.08}$
(iii) Consistyle (no K injection) $0.45\pm 0.12$ $32.73\pm 1.59$ $35.94\pm 1.71$ $0.32\pm 0.04$ $1.23\pm 0.25$ $0.84\pm 0.08$
(iv) Consistyle (no QK injection) $0.48\pm 0.11$ $32.60\pm 1.66$ $35.80\pm 1.65$ $\mathbf{0.18\pm 0.03}$ $\mathbf{0.72\pm 0.11}$ $\mathbf{0.94\pm 0.05}$
(v) Consistyle (w/o DIFT) $0.43\pm 0.12$ $\mathbf{32.88\pm 1.52}$ $35.74\pm 1.57$ $0.49\pm 0.03$ $2.19\pm 1.41$ $0.74\pm 0.14$
(vi) Consistyle (no AdaIN) $0.42\pm 0.12$ $32.81\pm 1.59$ $35.66\pm 1.71$ $0.23\pm 0.03$ $1.39\pm 0.27$ $0.82\pm 0.09$
Figure 6. Demonstration of the method’s limitations. The first row illustrates inconsistencies in generating a complex object (spaceship), where high visual detail leads to variation across images. The second row highlights a failure to align with a distinct style-specifically, the Papercraft Collage style, evident in the face details.
@@ 5. Limitations
Our approach has several limitations, as can be seen in Fig. 6. First, similar to Consistory (Tewel et al., 2024), it can produce suboptimal results when the correspondence mappings or cross-attention masks fail to accurately capture the intended relationships between images. Second, consistency can degrade for subjects with complex structures, such as large vehicles like boats and spaceships, where fine details are often challenging to maintain. This effect can extend to human subjects, where intricate facial features or clothing elements may vary across generations. The phenomenon is more prevalent when the initial subject interpretations differ significantly between images, such as generating a traditional wooden ship in one image and a modern yacht in another for the same ”boat” prompt. While this issue can sometimes be mitigated by selecting different seeds, it remains a potential weakness in cases where precise subject alignment is critical. Finally, highly distinctive artistic styles, such as Papercraft Collage, Voxel Art or other niche 3D aesthetics, can occasionally lead to style misalignment, as these styles often introduce unique structural deformations or exaggerated textures that challenge the statistical alignment technique we employ.
We also observed artifacts in some generated images. To verify that these do not originate from our method, we manually labeled 400 samples as having either minor artifacts (e.g., small visual glitches) or severe ones (e.g., extra limbs). We found that minor artifacts occurred in $18.8\%$ of SDXL images, $20.0\%$ of ours, and $23.2\%$ of Consistory; major artifacts appeared in $0\%, 2.2$$2\%$, and $4.0\%$ respectively, which indicates that this stems from the base model.
@@ 6. Conclusion
We present Consistyle, a training-free approach for improving consistency while preserving style alignment in text-to-image generation. Our method leverages attention manipulation to enable controlled characteristic sharing between images, even in cases with significant appearance differences. This demonstrates that consistent image generation is feasible in style-diverse contexts, despite the typical entanglement of style and content that often challenges.
Acknowledgements.
This work was supported by a grant from the Tel Aviv University Center for AI and Data Science (TAD). We would like to thank Yoad Tewel for his insightful feedback.
APPENDIX
Appendix A Training Details for DB LoRA and B4M
In this section, we provide a detailed overview of the training procedures used for the baselines we trained—DreamBooth LoRA (DB LoRA) and Break for Make (B4M).
A.1. DreamBooth LoRA
Workflow.
Instance images were generated with the first prompt of the prompt group, combined with the appropriate style from the style group. It was paired with class images generated with random seeds on the concept token to improve results while mitigating overfitting. Instance prompts were constructed by substituting the concept token with a unique placeholder token.
Parameter Details.
Models were trained for $250$ steps at a resolution of $1024\times 1024$, using a batch size of $1$ with gradient accumulation over $3$ steps. Mixed precision (fp16) and gradient checkpointing were applied to reduce memory usage. Optimization employed 8-bit Adam with a constant learning rate of $1\times 10^{-4}$ and no warmup. SNR weighting was applied with $\gamma=5.0$, and seeds were clamped to the 32-bit integer range for reproducibility.
A.2. Break for Make
The training procedure for B4M consists of three phases:
Phase 1: Content LoRA
We first train a LoRA for the content reference (e.g., ”a kitten”). For stable results, it is recommended to use at least three images. To this end, we generate one image with vanilla SDXL and two additional images with IP-Adapter in a ”photo-realistic” style. These prompts are deliberately different from those used in our evaluation templates. To disentangle content from style, we additionally provide three style reference images, as suggested by the original authors, chosen from style categories not included in our evaluation set. Training is performed for 1000 steps, as recommended. However, we observe that overfitting may occur; in some cases, reducing the training to 500 steps yields better results.
Phase 2: Style LoRA
Next, we train a LoRA for style references. We generate three images (e.g., landscape, cat, etc.) with a specific style description (e.g., pixel art, line-art). As in Phase 1, we train for 1000 steps and provide three additional content images to encourage disentanglement between content and style.
Phase 3: Content–Style LoRAs
In the final phase, we train separate LoRAs for each content–style pair, again for 1000 steps.
Inference
At inference time, images are generated by composing the learned tokens corresponding to content and style. For example, a prompt could be: *“snq woman coding on a laptop, w@z pixel art style.”*
Appendix B User Study Instructions and Questions